# Perturbation theory for gravitational shadows in Kerr-like spacetimes

### v1

@@ IV.1 Kerr-Newmann Black Hole
The components of the Kerr-Newman metric in the Benenti form Papadopoulos:2018nvd are given by
@@
To illustrate the accuracy of the derived approximation, we first recall the exact parametric expression for the Kerr-Newman shadow Kobialko:2022ozq; Kobialko:2023qzo:
@@
The primary advantage of having explicit analytical expressions for the shadow boundary is the significant simplification they provide for the inverse problem of extracting metric parameters from an observed shadow. In a general $r$-parametric approach, one must repeatedly map the orbital radius $r$ to the celestial angle $\Phi$ and shadow radius $R$ for each new set of trial parameters, as the mapping $r\rightarrow(\Phi,R)$ is parameter-dependent. Reconstructing parameters would then require constructing a large, pre-computed mapping or training a neural network over a broad parameter space—a computationally expensive process. In contrast, with an explicit expression $R(\Phi)$ available, one can directly formulate the fitting procedure as minimizing a loss function, e.g., $\sum\left(R_{\text{obs}}(\Phi_{i})-R_{\text{model}}(\Phi_{i})\right)^{2}$, using a standard optimizer such as Adam Kingma:2014vow. Leveraging this, we have developed a parameter estimation algorithm that employs our fifth-order analytical approximation to reconstruct the spin $a$ and the dimensionless charge $q$ of a Kerr-Newman black hole. As synthetic observational data, we used the exact parametric expressions for the shadow boundary. Predicted metric parameters for a fixed observer inclination $\bar{\theta}=\pi/2$ are presented in Table 1, demonstrating satisfactory accuracy. However, for other inclination angles $\bar{\theta}$, the results are less promising, as a higher-order perturbative expansion would be required to maintain accuracy.
@@ IV.2 Kerr-Newmann Black Hole in a plasma medium
The key requirement for shadow spectroscopy Pantig:2025deu is a frequency-dependent shift in the observed gravitational shadow. This condition is naturally satisfied in two scenarios: observations involving beams of massive particles Kobialko:2024zhc, or radiation propagating through a plasma medium Perlick:2015vta. In this work, we focus on and demonstrate the applicability of our analytical results to the latter case.
The components of Kerr-Newmann metric in the Benenti form taking into account the power law of plasma distribution are
@@
A key feature of these analytical formulas is the explicit dependence of the shadow morphology on the observational frequency $\omega$. For instance, from the behavior of $\delta C_{A}$ alone, one can directly estimate parameters like $\alpha$ or $\beta$ based on the slope of $\delta C_{A}$ plotted against $\omega^{-2}$. This frequency dependence lays the foundation for shadow spectroscopy, which allows for the efficient disentanglement and extraction of individual metric parameters from multi-frequency observations Kobialko:2024zhc.
@@ IV.3 Modified Kerr and Kerr-Sen black holes
Finally, we apply our formalism to more general metric families. One such multi-parameter family encompasses both Modified Kerr and Kerr-Sen black holes Papadopoulos:2018nvd. The corresponding metric components in the Benenti form are given by
@@
It should be noted that the same optimization method used successfully to reconstruct the charge in Kerr-Newman metrics proves less effective for disentangling individual parameters in more general multi-parameter families. A similar limitation has been observed in static spacetimes Kobialko:2024zhc; Pantig:2025deu. In such cases, multi-frequency observations can significantly enhance the robustness of parameter recovery. We anticipate that the analytical formulas derived in this work will facilitate future investigations of this challenging problem.

### v2

@@ IV.1 Kerr-Newmann Black Hole
The components of the Kerr-Newman metric in the Benenti-Francavilla form Papadopoulos and Kokkotas (2018) are given by
@@
To illustrate the accuracy of the derived approximation, we first recall the exact parametric expression for the Kerr-Newman shadow Kobialko et al. (2022b, 2024):
@@
The primary advantage of having explicit analytical expressions for the shadow boundary is the significant simplification they provide for the inverse problem of extracting metric parameters from an observed shadow Banerjee et al. (2020); Tsukamoto et al. (2014). In a general $r$-parametric approach, one must repeatedly map the orbital radius $r$ to the celestial angle $\Phi$ and shadow radius $R$ for each new set of trial parameters, as the mapping $r\rightarrow(\Phi,R)$ is parameter-dependent. Reconstructing parameters would then require constructing a large, pre-computed mapping or training a neural network over a broad parameter space—a computationally expensive process. In contrast, with an explicit expression $R(\Phi)$ available, one can directly formulate the fitting procedure as minimizing a loss function, e.g., $\sum\left(R_{\text{obs}}(\Phi_{i})-R_{\text{model}}(\Phi_{i})\right)^{2}$, using a standard optimizer such as Adam Kingma and Ba (2014). Leveraging this, we have developed a parameter estimation algorithm that employs our fifth-order analytical approximation to reconstruct the spin $a$ and the dimensionless charge $q$ of a Kerr-Newman black hole. As synthetic observational data, we used the exact parametric expressions for the shadow boundary. Predicted metric parameters for a fixed observer inclination $\bar{\theta}=\pi/2$ are presented in Table 1, demonstrating satisfactory accuracy. However, for other inclination angles $\bar{\theta}$, the results are less promising, as a higher-order perturbative expansion would be required to maintain accuracy.
@@ IV.2 Kerr-Newmann Black Hole in a plasma medium
The key requirement for shadow spectroscopy Pantig and Övgün (2025) is a frequency-dependent shift in the observed gravitational shadow. This condition is naturally satisfied in two scenarios: observations involving beams of massive particles Kobialko and Gal’tsov (2025), or radiation propagating through a plasma medium Perlick et al. (2015). In this work, we focus on and demonstrate the applicability of our analytical results to the latter case.
The components of Kerr-Newmann metric in the Benenti-Francavilla form taking into account the power law of plasma distribution are
@@
A key feature of these analytical formulas is the explicit dependence of the shadow morphology on the observational frequency $\omega$. For instance, from the behavior of $\delta C_{A}$ alone, one can directly estimate parameters like $\alpha$ or $\beta$ based on the slope of $\delta C_{A}$ plotted against $\omega^{-2}$. This frequency dependence lays the foundation for shadow spectroscopy, which allows for the efficient disentanglement and extraction of individual metric parameters from multi-frequency observations Kobialko and Gal’tsov (2025).
@@ IV.3 Modified Kerr and Kerr-Sen black holes
Finally, we apply our formalism to more general metric families. One such multi-parameter family encompasses both Modified Kerr and Kerr-Sen black holes Papadopoulos and Kokkotas (2018). The corresponding metric components in the Benenti-Francavilla form are given by
@@
It should be noted that the same optimization method used successfully to reconstruct the charge in Kerr-Newman metrics proves less effective for disentangling individual parameters in more general multi-parameter families. A similar limitation has been observed in static spacetimes Kobialko and Gal’tsov (2025); Pantig and Övgün (2025). In such cases, multi-frequency observations can significantly enhance the robustness of parameter recovery. We anticipate that the analytical formulas derived in this work will facilitate future investigations of this challenging problem.